OpenStudy (anonymous):

which one of the following is a complex zero of P(x)=x^4-16

OpenStudy (anonymous):

\[x^4-16=(x^2+4)(x^2-4)=(x^2+4)(x+2)(x-2)\]

OpenStudy (anonymous):

the real zeros are \(2\) and \(-2\)

OpenStudy (anonymous):

the complex ones you get from solving \[x^2+4=0\] do you know how to do that?

OpenStudy (anonymous):

wouldn't x=2?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[x^2+4=0\\ x^2=-2\\ x=\pm\sqrt{-2}\] or \[x=\pm\sqrt{2}i\]

OpenStudy (anonymous):

it is not \(x=2\) because \(2^2+4=4+4=8\) not zero

OpenStudy (anonymous):

okay so what if my choices are only..

OpenStudy (anonymous):

i made a typo there second line should be \(x^2=-4\)

OpenStudy (anonymous):

\(-2i\) and \(2i\) are the zeros pick \(-2i\) who is jsara?

OpenStudy (anonymous):

okay that makes sense.... and i just typed in random letters but I asked myself that same question as i was sending it hahahaha

OpenStudy (anonymous):

if those letters are random i will have a beer now